OpenStudy (shamil98):

implicit differentiation.. x^2 +xy + y^2 = 1

OpenStudy (shamil98):

\[\huge \frac{ d }{ dx } x^2 + \frac{ d }{ dx } xy + \frac{ d }{ dx }y^2 = \frac{ d }{ dx }1\]

OpenStudy (shamil98):

I use the product rule for xy right?

OpenStudy (anonymous):

yep

OpenStudy (shamil98):

soo f'g + fg' = y + xy' \[\huge 2x + y + xy' + 2yy' = 0\]

OpenStudy (anonymous):

looks good so far...

OpenStudy (shamil98):

\[\huge y'(2y+x) = -y -2x\] \[\Huge y' =\frac{ -y-2x }{ 2y+x }\] ?

OpenStudy (anonymous):